OpenStudy (anonymous):

what is the square root of 75/13?

OpenStudy (anonymous):

5 3/13

OpenStudy (anonymous):

You forgot the square root sign over the fraction 3/13

OpenStudy (anonymous):

Oh god thanks!

OpenStudy (anonymous):

Would you write out the correct answer with the square root sign so rotted sees it too?

OpenStudy (anonymous):

D.w he cud do it himself =)

OpenStudy (anonymous):

Ok, I'll do it, but no medal from me. Rotted, you solve it like this. \[\sqrt{75/13}\] becomes \[\sqrt{25*3/13}\] Which becomes \[\sqrt{25}*\sqrt{3/13}\] Which becomes \[5\sqrt{3/13}\]